OpenStudy (anonymous):

a cylindrical water trough has a diameter of 6 feet and a height of 2 feet. it is being filled at the rate of 4 ft^2/min. how fast is the water level rising when the water is 1 foot deep?

ganeshie8 (ganeshie8):

\(V = \pi r^2 h = 9 \pi h\) differentiate both sides and plugin the given values (note that typo in in question for flow rate - it should be 4 ft^3/min)